OpenStudy (andijo76):

Essay; show all work. A simple dartboard has three areas… the main board has a radius of 10 inches, there is a circle with a radius of 7 inches, and the bullseye has a radius of 3 inches. What is the probability of a random dart landing inside the bullseye

OpenStudy (andijo76):

thhere is more

OpenStudy (andijo76):

Round to the nearest thousandth hint use

OpenStudy (anonymous):

it is the area of the bull's eye divided by the total area

OpenStudy (andijo76):

hint use 

OpenStudy (anonymous):

compute the area of the dartboard using the formula you just wrote then compute the area of the bulls eye as well divide the second by the first

OpenStudy (anonymous):

\[\frac{\pi\times 3^2}{\pi\times 10^2}=\frac{9}{100}=.09\]